# AuroraCap: Efficient, Performant Video Detailed Captioning and a New Benchmark

### v1

@@ 3.1.1 Video Collection and Processing
To ensure the reliability of the benchmark, it is crucial to maintain high video quality, balanced data distribution, and content complexity. Panda-70M Chen et al. (2024e) offers a high-resolution, open-domain YouTube video dataset with diverse one-minute clips across wildlife, cooking, sports, news, TV shows, gaming, and 3D rendering, ideal for studying complex real-world scenarios. Additionally, a large volume of aesthetically appealing videos from user-uploaded platforms like Mixkit Mixkit (2024), Pixabay Pixabay (2024), and Pexels Pexels (2024) provides scenic views and visually pleasing human activities with minimal transitions and simpler events. Ego4D Grauman et al. (2022) complements the video source by focusing on ego-centric human activities and auto-driving scenarios, ensuring comprehensive coverage of real-world scenes. To mitigate content homogeneity among these candidate videos and maintain diversity in the final dataset, inspired by ShareGPT4Video Chen et al. (2024d), we building VDC upon those various video sources. Note that the videos used in VDC construction are not included in the training data of AuroraCap. To ensure balanced data distribution, we allocate equal proportions of videos from Panda-70M Chen et al. (2024e), Ego4D Grauman et al. (2022), Mixkit Mixkit (2024), Pixabay Pixabay (2024), and Pexels Pexels (2024). We first split the video into clips and apply dense frame extraction.
@@
Panda-70M Chen et al. (2024e) 229 22.25% 5,714 24.95 7.18
Ego4D Grauman et al. (2022) 196 19.05% 10,935 55.79 19.46
Mixkit Mixkit (2024) 197 19.14% 3,261 16.55 6.58
Pixabay Pixabay (2024) 199 19.34% 4,748 23.86 8.99
Pexels Pexels (2024) 208 20.21% 4,343 20.88 7.99
@@
3.1.2 Structured Detailed Captions Construction Pipeline
We believe that a comprehensive detailed video caption benchmark should encompass various aspects, including main objects, camera movements, and background. However, most existing benchmarks Li et al. (2024b); Chen et al. (2024d) provide only a single caption for the entire video with less structured details. Therefore, we develop a structured detailed captions construction pipeline to generate extra detailed descriptions from various perspectives, significantly extending the length and enhancing the richness compared to previous benchmarks. Following Ju et al. (2024), the structured captions in VDC encompass not only short and detailed captions but also three additional categories: (1) main object caption, offering a comprehensive analysis of the primary subjects’ actions, attributes, interactions, and movements across frames, including variations in posture, expression, and speed; (2) background caption, providing detailed descriptions of the background, such as objects, location, weather, time, and dynamic elements; and (3) camera caption, which details the camera work, including shot types, angles, movements, transitions, and special effects.
To generate detailed, fine-grained, and accurate captions, we leverage GPT-4o to produce video descriptions. We utilize the dense video frames to obtain captions. We observed that generating all captions in a single conversation round often introduces hallucinations in the detailed captions. To address this, we design a hierarchical prompt strategy to efficiently obtain accurate structured captions and detailed captions in two conversation rounds: (1) structured captions generation and (2) detailed captions integration. In the first round, the prompt briefly introduces the differences between structured captions and uses the dense video frames as input to generate the short caption, main object caption, background caption, camera caption, and the detailed caption. In the second round, the generated captions serve as the reference. The second-round prompt guides GPT-4o to enhance the detailed caption based on the initial captions, ensuring consistency without introducing new entities or relations, and producing a vivid, engaging, and informative description. The whole prompt template for the structured detailed captions construction pipeline can be found in Appendix H. Finally, we conduct a manual review to correct captions with hallucinations and supplement omitted visual elements. The refined detail structured captions are then used as the ground truth for evaluation.
@@ 3.1.3 Comparison on Numerical Statistics
Based on the hierarchical scheme, VDC can capture a rich variety of details of the video and reduce hallucinations. The visual representation in Figure 2 demonstrates the video duration distribution of VDC. Over $87\%$ of the videos exhibit a duration ranging from 10K to 12K frames, while $1\%$ of videos extending beyond 60 seconds. Only $13\%$ of videos have duration less than 10 seconds. As illustrated in Table 1, the average length of detailed descriptions in VDC is significantly longer than in previous benchmarks. Figure 2 shows the length distribution of structured captions in VDC, with detailed captions averaging over 500 words. Appendix J present more statistics.
@@ 3.2.1 VDCscore: Evaluating Detailed Captions with LLMs
Evaluating video captions requires not only assessing the quality of the captions but also flexibly evaluating the alignment between the video and the caption. While metrics such as BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), and ROUGE-L Lin (2004) have been employed for caption evaluation, these metrics are predominantly designed for short captions and rely heavily on word-level frequency-based alignment. Given the advanced semantic understanding capabilities of large language models (LLMs), Video-ChatGPT Li et al. (2023b) proposes using LLM as an evaluation assistant to directly judge the correctness of the whole predicted captions and assign scores. However, as demonstrated in Table 6, our experiments indicate that when dealing with detailed captions, the direct application of LLM struggles to accurately distinguish the correctness of various predicted captions, fails to effectively evaluate the precision of detailed descriptions, and exhibits a tendency to assign disproportionately lower scores. Therefore, we introduce VDCscore, a novel quantitative metric that utilizes LLMs to evaluate the similarity between predicted and ground-truth detailed captions through a divide-and-conquer approach.
@@
The core idea of VDCscore is to decompose long detailed captions into multiple short question-answering pairs, avergae the evaluation of each pair as the final result. We elaborate the design of VDCscore in the following parts: (1) ground-truth question-answer pairs extraction, (2) responsed answers generation and (3) answers matching. As illustrated in Figure 3, we first employ GPT-4o to generate question-answer pairs from the detailed ground-truth caption. To ensure that the generated question-answer pairs capture as much information as possible from the original caption and facilitate accurate evaluation by LLMs, we constrain the number of generated pairs and impose specific guidelines: the questions must be open-ended, and the answers should be concise, and directly relevant to the questions. For a fair comparison and to mitigate potential variability arising from generating different question-answer pairs for the same caption, we pre-generate a standardized set of question-answer pairs for all captions in VDC, as depicted in Figure 3. The used prompt templates used along with additional examples, are provided in Appendix I.
@@
VDCscore subsequently analyzes the predicted captions by leveraging ground-truth question-answer pairs. We prompt GPT-4o to read the detailed predicted captions and generate answers based solely on these captions. To mitigate biases arising from discrepancies in the length between ground-truth and predicted answers, we also impose constraints ensuring that responses are limited to concise sentences or phrases. Consequently, for each pair of ground-truth and predicted captions, we obtain a set of <question, correct answer, predicted answer> triplets. Following Video-ChatGPT Li et al. (2023b), we then ask GPT-4o to output two scores for each triplet: one for answer correctness and another for answer quality. The final accuracy and score are calculated by averaging the correctness score and quality score respectively. When using two same captions as input, VDCscore returns an accuracy of 100%, demonstrating the feasibility and reliability.
Figure 4: Comparison between various models with different number of visual tokens input on VDC. For Gemini-1.5-Pro, we only report the performance. We manage the number of visual tokens by managing token merging for AuroraCap, and manage the number of frames for others. AuroraCap achieves a much better VDCscore than all other models given a certain compression in the number of visual tokens kept and indeed approaches the performance of Gemini-1.5-Pro.
Table 6: Comparison of AuroraCap with LLM-based baseline methods on VDCscore and other evaluation metrics under zero-shot setting. For each evaluation metric, we report the average value of the five structured captions in VDC. Note that VDD, CIDEr, and BELU are only the average of background and main object caption, since the values of the others are closed to zero.
Model VDCscore VDD C B@1 B@4 M R Elo
Acc Score Acc Score
Gemini-1.5 Pro Reid et al. (2024) 41.73 2.15 49.68 3.07 5.97 29.72 2.63 21.21 20.19 1,171
LLaMA-VID Li et al. (2023f) 30.86 1.62 4.63 1.63 1.48 17.74 1.46 8.07 17.47 859
Video-ChatGPT-7B Maaz et al. (2023) 31.12 1.62 8.57 1.84 2.92 17.31 2.19 11.57 16.96 944
MovieChat-7B Song et al. (2023) 31.92 1.64 10.24 1.86 5.14 14.33 3.17 13.60 14.98 890
VILA-7B Lin et al. (2023b) 32.61 1.70 16.27 2.02 8.20 19.13 2.11 5.62 16.63 1,073
Video-LLaVA-7B Lin et al. (2023a) 32.80 1.72 14.14 2.00 4.43 17.20 2.32 10.36 17.53 1,007
LLaVA-1.5-7B Liu et al. (2023) 33.98 1.76 26.71 2.33 6.63 29.80 2.54 22.79 20.36 825
LongVA-7B Zhang et al. (2024b) 34.50 1.79 32.65 2.69 4.83 18.75 2.16 13.43 14.84 969
LLaVA-1.5-13B Liu et al. (2023) 34.78 1.80 28.26 2.36 3.90 20.43 2.02 26.37 17.87 943
LLaVA-NeXT-V7B Zhang et al. (2024d) 35.46 1.85 25.62 2.34 2.66 20.18 2.33 28.17 17.51 1,022
LLaVA-1.6-7B Liu et al. (2024a) 35.70 1.85 40.16 2.69 3.09 17.36 1.59 24.23 17.08 846
LLaVA-1.6-13B Liu et al. (2024a) 35.85 1.85 34.55 2.51 5.55 29.23 2.50 20.26 19.96 728
ShareGPT4Video-8B Chen et al. (2024d) 36.17 1.85 36.44 1.85 1.02 12.61 0.79 8.33 16.31 1,102
LLaVA-OV-7B Li et al. (2024a) 37.45 1.94 41.83 2.70 4.09 28.34 2.84 23.98 19.59 1,155
InternVL-2-8B Chen et al. (2023b) 37.72 1.96 48.99 3.03 5.59 15.75 2.48 10.76 17.63 1,081
AuroraCap-7B 38.21 1.98 48.33 2.90 9.51 30.90 4.06 19.09 21.58 1,267
@@
Table 7: Comparison of AuroraCap with LLM-based baseline methods on VDCscore under zero-shot structured captions setting. We consider the VDCscore of detailed captions, short captions, background captions, main object captions and camera captions. We also test the vision-blind case suggesting by Chen et al. (2024c); Tong et al. (2024a).
@@
Vicuna-v1.5-7B Chiang et al. (2023) 21.68 / 1.12 23.06 / 1.17 22.02 / 1.15 22.64 / 1.16 23.09 / 1.20
Llama-3.1-8B Dubey et al. (2024) 17.83 / 1.00 17.90 / 1.02 19.52 / 1.10 19.57 / 1.10 20.10 / 1.22
Gemini-1.5 Pro Reid et al. (2024) 38.68 / 2.05 35.71 / 1.85 43.84 / 2.23 47.32 / 2.41 43.11 / 2.22
LLaMA-VID Li et al. (2023f) 39.47 / 2.10 29.92 / 1.56 28.01 / 1.45 31.24 / 1.59 25.67 / 1.38
Video-ChatGPT-7B Maaz et al. (2023) 37.46 / 2.00 29.36 / 1.56 33.68 / 1.70 30.47 / 1.60 24.61 / 1.26
MovieChat-7B Song et al. (2023) 37.25 / 1.98 32.55 / 1.59 28.99 / 1.54 31.97 / 1.64 28.82 / 1.46
VILA-7B Lin et al. (2023b) 34.33 / 1.83 30.40 / 1.55 35.15 / 1.80 33.38 / 1.72 29.78 / 1.58
Video-LLaVA-7B Lin et al. (2023a) 37.48 / 1.97 30.67 / 1.63 32.50 / 1.70 36.01 / 1.85 27.36 / 1.43
LLaVA-1.5-7B Liu et al. (2023) 38.38 / 2.04 28.61 / 1.51 34.86 / 1.79 34.62 / 1.76 33.43 / 1.73
LongVA-7B Zhang et al. (2024b) 35.32 / 1.90 31.94 / 1.63 36.39 / 1.85 40.95 / 2.11 27.91 / 1.48
LLaVA-1.5-13B Liu et al. (2023) 38.97 / 2.07 30.89 / 1.60 34.79 / 1.78 36.27 / 1.84 33.00 / 1.74
LLaVA-NeXT-V7B Zhang et al. (2024d) 39.73 / 2.10 30.63 / 1.60 36.54 / 1.88 36.54 / 1.88 33.84 / 1.77
LLaVA-1.6-7B Liu et al. (2024a) 36.50 / 1.93 31.91 / 1.65 37.58 / 1.92 36.03 / 1.85 36.47 / 1.89
LLaVA-1.6-13B Liu et al. (2024a) 35.61 / 1.86 31.90 / 1.66 38.90 / 1.99 36.65 / 1.87 36.18 / 1.89
ShareGPT4Video-8B Chen et al. (2024d) 33.28 / 1.76 39.08 / 1.94 35.77 / 1.81 37.12 / 1.89 35.62 / 1.84
LLaVA-OV-7B Li et al. (2024a) 37.82 / 2.02 32.58 / 1.70 37.43 / 1.92 38.21 / 1.96 41.20 / 2.13
InternVL-2-8B Chen et al. (2023b) 39.08 / 2.11 33.02 / 1.74 37.47 / 1.89 44.16 / 2.22 34.89 / 1.82
Table 8: Comparison of AuroraCap with LLM-based baseline methods on VDD Liu et al. (2024a) under zero-shot structured captions setting. We consider the VDD Liu et al. (2024a) of detailed captions, short captions, background captions, main object captions and camera captions.
Model Camera Short Background Main Object Detailed
 Acc / Score  Acc / Score  Acc / Score  Acc / Score  Acc / Score
Gemini-1.5 Pro Reid et al. (2024) 18.89 / 2.115 16.91 / 1.572 57.73 / 3.263 41.64 / 2.886 2.581 / 0.330
LLaMA-VID Li et al. (2023f) 28.28 / 2.513 1.034 / 1.042 6.198 / 1.895 3.063 / 1.366 2.046 / 0.304
Video-ChatGPT-7B Maaz et al. (2023) 16.00 / 2.175 4.173 / 1.032 14.14 / 2.273 3.001 / 1.423 1.038 / 0.192
VILA-7B Lin et al. (2023b) 4.005 / 1.751 2.087 / 0.233 22.45 / 2.385 10.10 / 1.672 1.015 / 0.262
Video-LLaVA-7B Lin et al. (2023a) 20.00 / 2.336 3.193 / 1.064 17.17 / 2.253 11.11 / 1.765 3.130 / 0.316
LLaVA-1.5-7B Liu et al. (2023) 26.88 / 2.515 1.222 / 0.793 36.50 / 2.725 16.92 / 1.937 0.694 / 0.276
LongVA-7B Zhang et al. (2024b) 17.00 / 2.204 1.016 / 0.794 50.00 / 3.203 15.31 / 2.196 0.002 / 0.247
LLaVA-1.5-13B Liu et al. (2023) 32.65 / 2.662 2.836 / 0.922 37.55 / 2.749 18.98 / 1.978 0.688 / 0.275
LLaVA-NeXT-V7B Zhang et al. (2024d) 29.81 / 2.645 1.913 / 0.957 33.23 / 2.692 18.02 / 1.999 0.887 / 0.279
LLaVA-1.6-7B Liu et al. (2024a) 21.11 / 2.229 8.696 / 1.146 53.31 / 3.105 27.01 / 2.286 1.282 / 0.279
LLaVA-1.6-13B Liu et al. (2024a) 21.56 / 2.199 9.798 / 1.206 39.37 / 2.692 29.73 / 2.329 1.287 / 0.271
ShareGPT4Video-8B Chen et al. (2024d) 33.28 / 1.768 4.908 / 0.986 35.77 / 1.813 37.12 / 1.899 3.213 / 0.752
LLaVA-OV-7B Li et al. (2024a) 17.11 / 2.086 11.15 / 1.277 55.82 / 3.149 27.84 / 2.258 1.372 / 0.249
InternVL-2-8B Chen et al. (2023b) 29.00 / 2.545 4.041 / 1.079 63.64 / 3.446 34.34 / 2.627 3.032 / 0.394
@@
3.2.2 Benchmarking Video Detailed Captioning
To our knowledge, no standard evaluation benchmarks have been established for detailed video captioning. To advance this field, we assess several baselines on our proposed VDC. As illustrated in Table 6, we present a quantitative comparison between our AuroraCap with existing state-of-the-art LMMs. We compare the VDCscore with both rule-based and model-based caption metrics with AuroraCap performing well. BLEU Papineni et al. (2002), CIDEr Vedantam et al. (2015), ROUGE-L Lin (2004) and METEOR Banerjee & Lavie (2005) are included as representative rule-based metrics. For model-based metrics, we also consider Video Detailed Description Li et al. (2023b) (VDD), which employs ChatGPT as an evaluation assistant to compare full captions.
Table 7 presents VDCscore performance across various sections of structured captions within VDC. Following Chen et al. (2024c), we also incorporate vision-blind baselines. Furthermore, Figure 4 illustrates a schematic diagram of the performance and efficiency of different video training models. Since the comparison of model inference time under different architectures, models, deployment frameworks, and output lengths is unfair, so we used the number of visual tokens as a representation of efficiency. AuroraCap achieves superior performance in video detailed captioning while utilizing significantly fewer visual tokens than other models, fully highlighting the efficiency of AuroraCap. We also show additional experimental results with VDD metric as shown in Table 8. We also perform a human study using Elo ranking to supplement our evaluation and provide a more intuitive assessment of AuroraCap’s performance. As depicted in Figure 5, VDCscore shows the better correlation with human evaluation results than VDD and ROUGE metric.
@@ 3.2.2 Benchmarking Video Detailed Captioning
Figure 5: Pearson correlation analysis among three evaluation metrics— VDCscore, ROUGE-L Lin (2004), VDD Liu et al. (2024a), and human Elo rankings for video models. VDD reflects the score for detailed captions. VDCscore demonstrates the highest consistency with expert judgments, thereby reinforcing the reliability. The detailed settings are provided in the Appendix L.

### v2

@@ 3.1.1 Video Collection and Processing
To ensure the reliability of the benchmark, it is crucial to maintain high video quality, balanced data distribution, and content complexity. Panda-70M [20] offers a high-resolution, open-domain YouTube video dataset with diverse one-minute clips across wildlife, cooking, sports, news, TV shows, gaming, and 3D rendering, ideal for studying complex real-world scenarios. Additionally, a large volume of aesthetically appealing videos from user-uploaded platforms like Mixkit [95], Pixabay [101], and Pexels [100] provides scenic views and visually pleasing human activities with minimal transitions and simpler events. Ego4D [40] complements the video source by focusing on ego-centric human activities and auto-driving scenarios, ensuring comprehensive coverage of real-world scenes. To mitigate content homogeneity among these candidate videos and maintain diversity in the final dataset, inspired by ShareGPT4Video [19], we building VDC upon those various video sources. Note that the videos used in VDC construction are not included in the training data of AuroraCap. To ensure balanced data distribution, we allocate equal proportions of videos from Panda-70M [20], Ego4D [40], Mixkit [95], Pixabay [101], and Pexels [100] as shown in Table 5. We first split the video into clips and apply dense frame extraction.
@@
Panda-70M [20] 229 22.25% 5,714 24.95 7.18
Ego4D [40] 196 19.05% 10,935 55.79 19.46
Mixkit [95] 197 19.14% 3,261 16.55 6.58
Pixabay [101] 199 19.34% 4,748 23.86 8.99
Pexels [100] 208 20.21% 4,343 20.88 7.99
@@ 3.1.2 Structured Detailed Captions Construction Pipeline
We believe that a comprehensive detailed video caption benchmark should encompass various aspects, including main objects, camera movements, and background. However, most existing benchmarks [61, 19] provide only a single caption for the entire video with less structured details. Therefore, we develop a structured detailed captions construction pipeline to generate extra detailed descriptions from various perspectives, significantly extending the length and enhancing the richness compared to previous benchmarks. Following [51], the structured captions in VDC encompass not only short and detailed captions but also three additional categories: (1) main object caption, offering a comprehensive analysis of the primary subjects’ actions, attributes, interactions, and movements across frames, including variations in posture, expression, and speed; (2) background caption, providing detailed descriptions of the background, such as objects, location, weather, time, and dynamic elements; and (3) camera caption, which details the camera work, including shot types, angles, movements, transitions, and special effects.
To generate detailed, fine-grained, and accurate captions, we leverage GPT-4o to produce video descriptions. We utilize the dense video frames to obtain captions. We observed that generating all captions in a single conversation round often introduces hallucinations in the detailed captions. To address this, we design a hierarchical prompt strategy to efficiently obtain accurate structured captions and detailed captions in two conversation rounds: (1) structured captions generation and (2) detailed captions integration. In the first round, the prompt briefly introduces the differences between structured captions and uses the dense video frames as input to generate the short caption, main object caption, background caption, camera caption, and the detailed caption. In the second round, the generated captions serve as the reference. The second-round prompt guides GPT-4o to enhance the detailed caption based on the initial captions, ensuring consistency without introducing new entities or relations, and producing a vivid, engaging, and informative description. The whole prompt template for the structured detailed captions construction pipeline can be found in Appendix 15. Finally, we conduct a manual review to correct captions with hallucinations and supplement omitted visual elements. The refined detail structured captions are then used as the ground truth for evaluation.
@@ 3.1.3 Comparison on Numerical Statistics
Table 6: Comparison of AuroraCap with LLM-based baseline methods on VDCscore and other evaluation metrics under zero-shot setting. For each evaluation metric, we report the average value of the five structured captions in VDC. Note that VDD, CIDEr, and BLEU are only the average of background and main object caption, since the values of the others are closed to zero.
Model VDCscore VDD CI- BLEU BLEU Met- ROU- Elo
Acc Score Acc Score DEr @1 @4 eor GE
Gemini-1.5 Pro [105] 41.73 2.15 49.68 3.07 5.97 29.72 2.63 21.21 20.19 1,171
LLaMA-VID [72] 30.86 1.62 4.63 1.63 1.48 17.74 1.46 8.07 17.47 859
Video-ChatGPT-7B [87] 31.12 1.62 8.57 1.84 2.92 17.31 2.19 11.57 16.96 944
MovieChat-7B [114] 31.92 1.64 10.24 1.86 5.14 14.33 3.17 13.60 14.98 890
VILA-7B [77] 32.61 1.70 16.27 2.02 8.20 19.13 2.11 5.62 16.63 1,073
Video-LLaVA-7B [75] 32.80 1.72 14.14 2.00 4.43 17.20 2.32 10.36 17.53 1,007
LLaVA-1.5-7B [80] 33.98 1.76 26.71 2.33 6.63 29.80 2.54 22.79 20.36 825
LongVA-7B [157] 34.50 1.79 32.65 2.69 4.83 18.75 2.16 13.43 14.84 969
LLaVA-1.5-13B [80] 34.78 1.80 28.26 2.36 3.90 20.43 2.02 26.37 17.87 943
LLaVA-NeXT-V7B [161] 35.46 1.85 25.62 2.34 2.66 20.18 2.33 28.17 17.51 1,022
LLaVA-1.6-7B [81] 35.70 1.85 40.16 2.69 3.09 17.36 1.59 24.23 17.08 846
LLaVA-1.6-13B [81] 35.85 1.85 34.55 2.51 5.55 29.23 2.50 20.26 19.96 728
ShareGPT4Video-8B [19] 36.17 1.85 36.44 1.85 1.02 12.61 0.79 8.33 16.31 1,102
LLaVA-OV-7B [60] 37.45 1.94 41.83 2.70 4.09 28.34 2.84 23.98 19.59 1,155
InternVL-2-8B [23] 37.72 1.96 48.99 3.03 5.59 15.75 2.48 10.76 17.63 1,081
AuroraCap-7B 38.21 1.98 48.33 2.90 9.51 30.90 4.06 19.09 21.58 1,267
Based on the hierarchical scheme, VDC can capture a rich variety of details of the video and reduce hallucinations. The visual representation in Figure 2 demonstrates the video duration distribution of VDC. Over $87\%$ of the videos exhibit a duration ranging from 10K to 12K frames, while $1\%$ of videos extending beyond 60 seconds. Only $13\%$ of videos have duration less than 10 seconds. As illustrated in Table 1, the average length of detailed descriptions in VDC is significantly longer than in previous benchmarks. Figure 2 shows the length distribution of structured captions in VDC, with detailed captions averaging over 500 words. Appendix 17 present more statistics.
Table 7: Comparison of AuroraCap with LLM-based baseline methods on VDCscore under zero-shot structured captions setting. We consider the VDCscore of detailed captions, short captions, background captions, main object captions and camera captions. We also test the vision-blind case suggesting by [17, 124].
Model Camera Short Background Main Object Detailed
 Acc / Score  Acc / Score  Acc / Score  Acc / Score  Acc / Score
Vicuna-v1.5-7B [25] 21.68 / 1.12 23.06 / 1.17 22.02 / 1.15 22.64 / 1.16 23.09 / 1.20
Llama-3.1-8B [34] 17.83 / 1.00 17.90 / 1.02 19.52 / 1.10 19.57 / 1.10 20.10 / 1.22
Gemini-1.5 Pro [105] 38.68 / 2.05 35.71 / 1.85 43.84 / 2.23 47.32 / 2.41 43.11 / 2.22
LLaMA-VID [72] 39.47 / 2.10 29.92 / 1.56 28.01 / 1.45 31.24 / 1.59 25.67 / 1.38
Video-ChatGPT-7B [87] 37.46 / 2.00 29.36 / 1.56 33.68 / 1.70 30.47 / 1.60 24.61 / 1.26
MovieChat-7B [114] 37.25 / 1.98 32.55 / 1.59 28.99 / 1.54 31.97 / 1.64 28.82 / 1.46
VILA-7B [77] 34.33 / 1.83 30.40 / 1.55 35.15 / 1.80 33.38 / 1.72 29.78 / 1.58
Video-LLaVA-7B [75] 37.48 / 1.97 30.67 / 1.63 32.50 / 1.70 36.01 / 1.85 27.36 / 1.43
LLaVA-1.5-7B [80] 38.38 / 2.04 28.61 / 1.51 34.86 / 1.79 34.62 / 1.76 33.43 / 1.73
LongVA-7B [157] 35.32 / 1.90 31.94 / 1.63 36.39 / 1.85 40.95 / 2.11 27.91 / 1.48
LLaVA-1.5-13B [80] 38.97 / 2.07 30.89 / 1.60 34.79 / 1.78 36.27 / 1.84 33.00 / 1.74
LLaVA-NeXT-V7B [161] 39.73 / 2.10 30.63 / 1.60 36.54 / 1.88 36.54 / 1.88 33.84 / 1.77
LLaVA-1.6-7B [81] 36.50 / 1.93 31.91 / 1.65 37.58 / 1.92 36.03 / 1.85 36.47 / 1.89
LLaVA-1.6-13B [81] 35.61 / 1.86 31.90 / 1.66 38.90 / 1.99 36.65 / 1.87 36.18 / 1.89
ShareGPT4Video-8B [19] 33.28 / 1.76 39.08 / 1.94 35.77 / 1.81 37.12 / 1.89 35.62 / 1.84
LLaVA-OV-7B [60] 37.82 / 2.02 32.58 / 1.70 37.43 / 1.92 38.21 / 1.96 41.20 / 2.13
InternVL-2-8B [23] 39.08 / 2.11 33.02 / 1.74 37.47 / 1.89 44.16 / 2.22 34.89 / 1.82
Figure 4: Comparison between various models with different number of visual tokens input on VDC. For Gemini-1.5-Pro, we only report the performance. We manage the number of visual tokens by managing token merging for AuroraCap, and manage the number of frames for others. AuroraCap achieves a much better VDCscore than all other models given a certain compression in the number of visual tokens and approaches the SoTA performance.
@@ 3.2.1 VDCscore: Evaluating Detailed Captions with LLMs
Evaluating video captions requires not only assessing the quality of the captions but also flexibly evaluating the alignment between the video and the caption. While metrics such as BLEU [99], CIDEr [127], and ROUGE-L [76] have been employed for caption evaluation, these metrics are predominantly designed for short captions and rely heavily on word-level frequency-based alignment. Given the advanced semantic understanding capabilities of large language models (LLMs), Video-ChatGPT [64] proposes using LLM as an evaluation assistant to directly judge the correctness of the whole predicted captions and assign scores. However, as demonstrated in Table 6, our experiments indicate that when dealing with detailed captions, the direct application of LLM struggles to accurately distinguish the correctness of various predicted captions, fails to effectively evaluate the precision of detailed descriptions, and exhibits a tendency to assign disproportionately lower scores. Therefore, we introduce VDCscore, a novel quantitative metric that utilizes LLMs to evaluate the similarity between predicted and ground-truth detailed captions through a divide-and-conquer approach.
@@
The core idea of VDCscore is to decompose long detailed captions into multiple short question-answering pairs, avergae the evaluation of each pair as the final result. We elaborate the design of VDCscore in the following parts: (1) ground-truth question-answer pairs extraction, (2) responsed answers generation and (3) answers matching. As illustrated in Figure 3, we first employ GPT-4o***GPT-4o latest version in September, 2024. to generate question-answer pairs from the detailed ground-truth caption. To ensure that the generated question-answer pairs capture as much information as possible from the original caption and facilitate accurate evaluation by LLMs, we constrain the number of generated pairs and impose specific guidelines: the questions must be open-ended, and the answers should be concise, and directly relevant to the questions. For a fair comparison and to mitigate potential variability arising from generating different question-answer pairs for the same caption, we pre-generate a standardized set of question-answer pairs for all captions in VDC, as depicted in Figure 3. The used prompt templates used along with additional examples, are provided in Appendix 16.
@@
VDCscore subsequently analyzes the predicted captions by leveraging ground-truth question-answer pairs. We prompt GPT-4o to read the detailed predicted captions and generate answers based solely on these captions. To mitigate biases arising from discrepancies in the length between ground-truth and predicted answers, we also impose constraints ensuring that responses are limited to concise sentences or phrases. Consequently, for each pair of ground-truth and predicted captions, we obtain a set of <question, correct answer, predicted answer> triplets. Following Video-ChatGPT [64], we then ask GPT-4o to output two scores for each triplet: one for answer correctness and another for answer quality. The final accuracy and score are calculated by averaging the correctness score and quality score respectively. When using two same captions as input, VDCscore returns an accuracy of 100%, demonstrating the feasibility and reliability.
3.2.2 Benchmarking Video Detailed Captioning
To our knowledge, no standard evaluation benchmarks have been established for detailed video captioning. To advance this field, we assess several baselines on our proposed VDC. As illustrated in Table 6, we present a quantitative comparison between our AuroraCap with existing state-of-the-art LMMs. We compare the VDCscore with both rule-based and model-based caption metrics with AuroraCap performing well. BLEU [99], CIDEr [127], ROUGE-L [76] and METEOR [6] are included as representative rule-based metrics. For model-based metrics, we also consider Video Detailed Description [64] (VDD), which employs ChatGPT as an evaluation assistant to compare.
@@ 3.2.2 Benchmarking Video Detailed Captioning
Table 8: Comparison of AuroraCap with LLM-based baseline methods on VDD [81] under zero-shot structured captions setting. We consider the VDD [81] of detailed captions, short captions, background captions, main object captions and camera captions.
@@
Gemini-1.5 Pro [105] 18.89 / 2.115 16.91 / 1.572 57.73 / 3.263 41.64 / 2.886 2.581 / 0.330
LLaMA-VID [72] 28.28 / 2.513 1.034 / 1.042 6.198 / 1.895 3.063 / 1.366 2.046 / 0.304
Video-ChatGPT-7B [87] 16.00 / 2.175 4.173 / 1.032 14.14 / 2.273 3.001 / 1.423 1.038 / 0.192
VILA-7B [77] 4.005 / 1.751 2.087 / 0.233 22.45 / 2.385 10.10 / 1.672 1.015 / 0.262
Video-LLaVA-7B [75] 20.00 / 2.336 3.193 / 1.064 17.17 / 2.253 11.11 / 1.765 3.130 / 0.316
LLaVA-1.5-7B [80] 26.88 / 2.515 1.222 / 0.793 36.50 / 2.725 16.92 / 1.937 0.694 / 0.276
LongVA-7B [157] 17.00 / 2.204 1.016 / 0.794 50.00 / 3.203 15.31 / 2.196 0.002 / 0.247
LLaVA-1.5-13B [80] 32.65 / 2.662 2.836 / 0.922 37.55 / 2.749 18.98 / 1.978 0.688 / 0.275
LLaVA-NeXT-V7B [161] 29.81 / 2.645 1.913 / 0.957 33.23 / 2.692 18.02 / 1.999 0.887 / 0.279
LLaVA-1.6-7B [81] 21.11 / 2.229 8.696 / 1.146 53.31 / 3.105 27.01 / 2.286 1.282 / 0.279
LLaVA-1.6-13B [81] 21.56 / 2.199 9.798 / 1.206 39.37 / 2.692 29.73 / 2.329 1.287 / 0.271
ShareGPT4Video-8B [19] 33.28 / 1.768 4.908 / 0.986 35.77 / 1.813 37.12 / 1.899 3.213 / 0.752
LLaVA-OV-7B [60] 17.11 / 2.086 11.15 / 1.277 55.82 / 3.149 27.84 / 2.258 1.372 / 0.249
InternVL-2-8B [23] 29.00 / 2.545 4.041 / 1.079 63.64 / 3.446 34.34 / 2.627 3.032 / 0.394
@@
Refer to caption Refer to caption Refer to caption
Figure 5: Pearson correlation analysis among three evaluation metrics—VDCscore, ROUGE-L [76], VDD [81]—and human Elo rankings for video models. VDD reflects the score for detailed captions. VDCscore demonstrates the highest consistency with expert judgments, thereby reinforcing the reliability. The detailed settings are provided in the Appendix 19.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Table 7 presents VDCscore performance across various sections of structured captions within VDC. Following [17], we also incorporate vision-blind baselines. Furthermore, Figure 4 illustrates a schematic diagram of the performance and efficiency of different video training models. Since the comparison of model inference time under different architectures, models, deployment frameworks, and output lengths is unfair, so we used the number of visual tokens as a representation of efficiency. AuroraCap achieves superior performance in video detailed captioning while utilizing significantly fewer visual tokens than other models, fully highlighting the efficiency of AuroraCap. We also show additional experimental results with VDD metric as shown in Table 8. We also perform a human study using Elo ranking to supplement our evaluation and provide a more intuitive assessment of AuroraCap’s performance. As depicted in Figure 5, VDCscore shows better correlation with human evaluation results than VDD and ROUGE metric.